the average score of the first six tests was 84.the first five scores were 92,84,79,87 and 73.what was the score of the sixth test
Answer by oscargut(2103)

We have to find x where:
(92+84+79+87+73+x)/6 = 84
x+415 = 6(84)
x+415 = 504
x = 89
Answer: 89
If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com